The total of three sisters' ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. How old are the sisters?
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

The total of three sisters' ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. How old are the sisters?
-----
Equations:
D + M + J = 39
D = (1/2)M
D = J - 3
-------------------
Modify:
D + M + J = 39
M = 2D
J = D + 3
------------------
Substitute for "M" and for "J": solve for "D":
D + 2D + D+3 = 39
4D = 36
D = 9 years (Dina's age now)
M = 2D = 18 years (Michelle's age now)
J = D+3 = 12 years (Juliette's age now)
===========================================
Cheers,
Stan H.
===============

Answer by lynnlo(4176)

D+M+J=39
D===============9
M=============18
J===========12